The number of mole of atoms in 8 mol of Ca3 (PO4)2
1
Expert's answer
2020-04-28T12:20:46-0400

Number of mole of atoms in "1" mole of "Ca_3 (PO_4)_2=3+2\times (1+4)=13"

Number of moles of atoms in "8" moles of "Ca_3(PO_4)_2=13\times 8=104"